OpenStudy (anonymous):

when you have a word problem, how do you determine the variables and how do you determine what is dependent and independent?

OpenStudy (anonymous):

usually what you are looking for will be represented by a (or the) dependent vairable.

OpenStudy (anonymous):

@pgpilot326 would you mind if I send you the problem ?

OpenStudy (anonymous):

try another of the heavy hitters... i've actually got to run right now.

OpenStudy (anonymous):

@pgpilot326 Ok . Thanks though !

OpenStudy (anonymous):

@satellite73 can you help this person?

OpenStudy (anonymous):

have a good day!

OpenStudy (anonymous):

@pgpilot326 Thank you, you do the same.

OpenStudy (anonymous):

sure what is the question exactly?

OpenStudy (anonymous):

@satellite73 Write the equation for each situation. Do not solve. Define variables and state which one is independent and which one is dependent. Jason spends $1,100 buying supplies and equipments for his cleaning business. He charges $12 per hour. Write an equation for the profit that he would earn for hours worked.

OpenStudy (anonymous):

if he works 100 hours he makes \(\$12\times 100=\$1200\) for a profit of \($1200-1100=\$100\)

OpenStudy (anonymous):

if he works 200 hours he makes \(\$12\times 200=\$2400\) for a profit of \(2400-1100=\$1300\)

OpenStudy (anonymous):

now lets say he works \(x\) hours

OpenStudy (anonymous):

replace \(100\) or \(200\) by \(x\) and see that he makes \(\$12\times x\) total, and therfore a profit of \(12x-1100\)

OpenStudy (anonymous):

ok, I see. Now how do we determine the variables and what is dependent and what is independent ?

OpenStudy (anonymous):

the "independent variable" is what i called \(x\) the number of hours worked

OpenStudy (anonymous):

the "dependent variable' i did not use that would be the amount of profit

OpenStudy (anonymous):

if we called the profit \(y\) then we could write \[y=12x-1100\]

OpenStudy (anonymous):

if you are a bit confused by the "independent dependent" business you should be in this case we are supposed to think that the number of hours worked is the dependent variable, because the amount of money earned "depends" on the number of hours worked

OpenStudy (anonymous):

@satellite73 ohhhh ok I get it, because I was a bit confused. This is what I have, it may be wrong. y=12x-1100 ?

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

@Yayyy !!! ok. Thank you so much ! I really do appreciate all of your help.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yw